OpenStudy (anonymous):

the greatest height to which a man can throw a stone is H. the greatest distance to which he can throw it will be?

OpenStudy (anonymous):

Let the velocity with which the stone is thrown is v. Initially K.E. = 1/2 m v^2 Finally P.E. =mgH Initially K.E = Final P.E 1/2 mv^2 = mgH v = sqrt(2gH) This is the maximum velocity with which man can throw. The maximum distance one can throw = v^2/2g (Read about projectile motion if you want to know about this formula) Put the value of v in the above equation and

OpenStudy (anonymous):

......find the maximum distance.

OpenStudy (missmob):

hmmm

OpenStudy (irishboy123):

to max out on distance, the guy must throw the ball at the optimal angle, which is 45deg to the ground. i can show where the 45deg comes from if required. this means that he throws the ball with a horizontal velocity of sqrt(2gH)*cos(45deg) . the time take is the same as the time for the ball to rise and fall in the vertical plane. that follows from v = u + at giving t = (v-u)/g. specifically, t = 2 * sqrt(2gH)sin45deg/g working through the distance * time figures gives 4H*sin45*cos45 = 2H.

OpenStudy (irishboy123):

the key to all of this is knowing the basic equations for motion at constant acceleration, and applying them correctly in the horizontal and vertical axes. v = u + at (1) x = ut + (1/2)at^2 (2) V^2 = u^2 + 2ax (3) these should just trip off the tongue. the other equations for energy follow these. eg (1/2)mv^2 follows from (3), and mgh from (2).